### v1

# Randomized Kaczmarz with geometrically smoothed momentum

Abstract.
This paper studies the effect of adding geometrically smoothed momentum to the randomized Kaczmarz algorithm, which is an instance of stochastic gradient descent on a linear least squares loss function. We prove a result about the expected error in the direction of singular vectors of the matrix defining the least squares loss. We present several numerical examples illustrating the utility of our result and pose several questions.
Key words and phrases:
Momentum, randomized Kaczmarz, stochastic gradient descent
2020 Mathematics Subject Classification:
Primary 65F10; Secondary 37N30, 60H25, 65F20
1. Introduction
1.1. Introduction
Let $A\in\mathbb{R}^{m\times n}$ be a tall ($m\geq n$) matrix, $b\in\mathbb{R}^{m}$ be a column vector, and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be the least squares loss function
(1) $f(x)=\frac{1}{2}\|Ax-b\|_{2}^{2}=\frac{1}{2}\sum_{i=1}^{m}(\langle a_{i},x\rangle-b_{i})^{2}=\sum_{i=1}^{m}f_{i}(x),$
[…]
(2) $x_{k+1}=x_{k}-\alpha_{k}\nabla\frac{1}{p_{i_{k}}}f_{i_{k}}(x)=x_{k}-\alpha_{k}\frac{\langle a_{i_{k}},x_{k}\rangle-b_{i_{k}}}{p_{i_{k}}}a_{i_{k}},$
where $\alpha_{k}$ is the learning rate, and $i_{k}$ is chosen independently and randomly from $\{1,\ldots,m\}$ such that each $i$ is chosen with probability $p_{i}$. Setting $p_{i}=\|a_{i}\|_{2}^{2}/\|A\|_{F}^{2}$ and $\alpha_{k}=1/\|A\|_{F}^{2}$, where $\|\cdot\|_{F}$ is the Frobenius norm, results in the randomized Kaczmarz algorithm
(3) $x_{k+1}=x_{k}+\frac{b_{i_{k}}-\langle a_{i_{k}},x_{k}\rangle}{\|a_{i_{k}}\|_{2}^{2}}a_{i_{k}},$
see [19]. The randomized Kaczmarz algorithm has a geometric interpretation: at iteration $k$, we perform an orthogonal projection of $x_{k}$ on the affine hyperplane $\{y\in\mathbb{R}^{n}:\langle a_{i_{k}},y\rangle=b_{i_{k}}\}$ defined by the $i_{k}$-th equation. If $Ax=b$ is a consistent linear system of equations, meaning that there exists a solution $x$ satisfying $\langle a_{i},x\rangle=b_{i}$ for all $i\in\{1,\ldots,m\}$, then the randomized Kaczmarz algorithm has the following convergence rate result established in 2009 by Strohmer and Vershynin: by [25, Theorem 2] we have
(4) $\mathbb{E}\|x_{k}-x\|_{2}^{2}\leq(1-\eta)^{k}\|x-x_{0}\|_{2}^{2},$
where $\eta:=\sigma_{n}^{2}/\|A\|_{F}^{2}$, where $\sigma_{n}$ denotes the smallest singular value of $A$. In 2021, Steinerberger [23] proved several results about how the error decays in different directions. Let $v_{l}$ be the right singular vector of $A$ associated with the $l$-th largest singular value $\sigma_{l}$ of $A$. Then by [23, Theorem 1.1] we have
(5) $\mathbb{E}\langle x_{k}-x,v_{l}\rangle=(1-\eta_{\ell})^{k}\langle x_{0}-x,v_{l}\rangle,$
where $\eta_{l}:=\sigma_{l}^{2}/\|A\|_{F}^{2}$. Moreover, the direction in which the vectors approach the solution remains roughly constant: By [23, Theorem 1.3], we have
(6) $\mathbb{E}\left\langle\frac{x_{k}-x}{\|x_{k}-x\|_{2}},\frac{x_{k+1}-x}{\|x_{k+1}-x\|_{2}}\right\rangle^{2}=1-\frac{1}{\|A\|_{F}^{2}}\left\|A\frac{x_{k}-x}{\|x_{k}-x\|_{2}}\right\|_{2}^{2},$
assuming that $x_{k},x_{k+1}\not=x$. Moreover, Steinerberger poses the following question based on these results:
“Once $x_{k}$ is mainly comprised of small singular vectors, its direction does not change very much anymore …Could this property be used for convergence acceleration?” [23, Page 4].
Our main result addresses this question: We show that adding geometrically smoothed momentum to the randomized Kaczmarz algorithm can accelerate convergence in the direction of small singular vectors. More precisely, we define the randomized Kaczmarz with geometrically smoothed momentum (KGSM) algorithm, see (9), and establish Theorem 1.1, which extends (5) to a setting with geometrically smoothed momentum. Our method is interpretable and illustrated by several numerical examples in §2.
Before going into the details of the algorithm and results, we present a numerical example to illustrate the accelerated convergence and interesting dynamics of KGSM (9) compared to the randomized Kaczmarz (3), see Figure 1.
Figure 1. The error $|\langle x_{k}-x,v_{n}\rangle|$ in the direction of the smallest singular vector $v_{n}$ for randomized Kaczmarz (3) and KGSM (9) (see §2.3 for a precise description of this numerical example).
1.2. Motivation
We are motivated by Polyak’s Heavy Ball Momentum [22], which is the iteration
(7) $x_{k+1}=x_{k}-\alpha_{k}\nabla f(x_{k})+\beta_{k}(x_{k}-x_{k-1}),$
where $f$ is the loss function; the name momentum comes from interpreting $\beta_{k}$ and $x_{k}-x_{k-1}$ as mass and velocity, respectively. One potential idea is to translate (7) directly into the following modified randomized Kaczmarz algorithm
(8) $\left\{\begin{array}[]{ccl}x_{k+1}&=&x_{k}+\displaystyle\frac{{b}_{i_{k}}-\langle x_{k},a_{i_{k}}\rangle}{||a_{i_{k}}||^{2}_{2}}a_{i_{k}}+My_{k}\\[10.0pt] y_{k+1}&=&x_{k+1}-x_{k},\end{array}\right.$
where $y_{k}$ represents velocity and $M$ represents mass. However, numerical and theoretical results indicate this iteration is not effective [1, 7]; roughly speaking, the issue is that the momentum term is too noisy. One potential way to address this issue is to consider a minibatch version of Heavy Ball Momentum, which is studied by Bollapragada, Chen, and Ward [1], who prove that the convergence is accelerated when the minibatch size is large enough but at an increased computational cost per iteration, where minibatch size refers to the number of equations used to form the stochastic gradient, see [1, Eq. 1.2].
In this paper, we take a different approach: instead of using a larger minibatch, we geometrically smooth the momentum term. For $\beta\in[0,1)$ and $M\in[0,1]$, we define Kaczmarz with geometrically smoothed momentum (KGSM) by
(9) $\left\{\begin{array}[]{ccl}x_{k+1}&=&x_{k}+\displaystyle\frac{{b}_{i_{k}}-\langle x_{k},a_{i_{k}}\rangle}{||a_{i_{k}}||^{2}_{2}}a_{i_{k}}+My_{k}\\[10.0pt] y_{k+1}&=&\beta y_{k}+(1-\beta)(x_{k+1}-x_{k}).\end{array}\right.$
Note that setting $\beta=0$ results in (8), while setting $\beta$ close to $1$ results in a high level of smoothing. Our main result is an analog of (5) which has an improved convergence rate for the signed error in the direction of small singular vectors; see Theorem 1.1. We demonstrate numerically that this expected signed error in the direction of the singular vectors is an effective model for the absolute numerical error in the direction of the singular vectors in some situations. Furthermore, we pose several questions about the dynamics of KGSM in §4.
1.3. Related work
The Kaczmarz algorithm first appeared in the literature in a 1937 paper of Kaczmarz [8]. It was rediscovered in 1970 in the context of electron microscopy under the name algebraic reconstruction techniques (ART) [3]. In 2009, Strohmer and Vershynin [25] established the linear convergence result (4) for the randomized Kaczmarz algorithm for consistent linear systems of equations; this analysis was extended to noisy linear systems by Needell [17]. Moreover, the block Kaczmarz methods corresponding to minibatch stochastic gradient descent have been considered by several authors, see [1, 2, 16, 18, 19, 20].
In 2023, Marshall and Mickelin [14], considered the problem of determining an optimal learning rate schedule $\alpha_{k}$ for the randomized Kaczmarz algorithm
$x_{k+1}=x_{k}+\alpha_{k}\frac{\tilde{b}_{i_{k}}-\langle a_{i_{k}},x_{k}\rangle}{\|a_{i_{k}}\|^{2}_{2}}a_{i_{k}},$
for the case where $Ax=b$ is a consistent linear system of equations, and $\tilde{b}=b+\varepsilon$, where $\varepsilon$ has independent mean zero random entries. The main result of [14] shows that there is an optimal learning rate schedule that depends on two hyperparameters. A similar learning rate problem for block Kaczmarz problems was subsequently studied in [26].
Part of the motivation for the current paper was to understand adaptive learning rates in the context of the randomized Kaczmarz algorithm. In particular, we were motivated by trying to understand ADAM [9], a popular stochastic optimization method that adaptively controls the learning rate. The geometric smoothing of the momentum in (9) is somewhat reminiscent of the geometric averaging used in ADAM. We note that the idea of using weighted averaging (without momentum) was considered in the context of the randomized Kaczmarz algorithm by [13].
Several authors have considered Kaczmarz algorithms that can be interpreted in terms of adaptive learning rates in the context of solving linear systems of equations with sparse corrupted equations [5, 12, 23, 24]. Here, the challenge is to find a way to ignore the corrupted equations in a hard or soft way, which can be interpreted as using an adaptive learning rate that de-emphasizes certain equations.
Another line of related research is the accelerated randomized Kaczmarz (ARK) [10], which applies Nesterov’s accelerated procedure, which defines $x_{k},y_{k},v_{k}$ iteratively by
$\left\{\begin{array}[]{ccl}y_{k}&=&\alpha_{k}v_{k}+(1-\alpha_{k})x_{k},\\[5.0pt] x_{k+1}&=&y_{k}-\theta_{k}\nabla f(y_{k}),\\[5.0pt] v_{k+1}&=&\beta_{k}v_{k}+(1-\beta_{k})y_{k}-\gamma_{k}\nabla f(y_{k}),\\[5.0pt] \end{array}\right.$
see [21]. Modifications and extensions of ARK have been considered by several authors; see [4, 28, 29]. While our motivation is different, the iteration considered in this paper has some similarities to Nesterov’s accelerated procedure. In the general setting, several authors have considered momentum in stochastic iteration in different contexts; in particular, see [11, 15].
Another related work by Han, Su, Xie [6], incorporates momentum into randomized Douglas-Rachford methods, which are related to the randomized Kaczmarz algorithm and use iterative reflections. In this work, the authors use a sequence of reflections followed by one momentum-based step, which can be viewed as another way to achieve smoothing of the momentum term.
1.4. Main result
Our main result establishes a convergence result for the signed error in the direction of singular vectors for randomized Kaczmarz with geometrically smoothed momentum (KGSM). Let $\beta\in[0,1)$ and $M\in[0,1]$ be given parameters, suppose that $x_{0}\in\mathbb{R}^{n}$ is a given initial vector, and let $y_{0}=0\in\mathbb{R}^{n}$ be the zero vector. Define
(10) $\left\{\begin{array}[]{ccl}x_{k+1}&=&x_{k}+\displaystyle\frac{{b}_{i_{k}}-\langle x_{k},a_{i_{k}}\rangle}{||a_{i_{k}}||^{2}_{2}}a_{i_{k}}+My_{k}\\[10.0pt] y_{k+1}&=&\beta y_{k}+(1-\beta)(x_{k+1}-x_{k}),\end{array}\right.$
for $k=0,1,\ldots$ where $i_{k}$ is chosen randomly and independently from $\{1,\ldots,m\}$ such that $i$ is chosen with probability $\|a_{i}\|_{2}^{2}/\|A\|_{F}^{2}$.
Theorem 1.1 (Main result).
Fix $\beta\in[0,1)$, $M\in[0,1]$, and $\ell\in\{1,\ldots,n\}$. Suppose that $x_{k}$ is defined by KGSM (10). For all $k\geq 0$ we have
(11) $\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k}\begin{bmatrix}1\\ \frac{-1}{1-\beta}\end{bmatrix}\langle x_{0}-x,v_{l}\rangle,$
where
[…]
The proof of Theorem 1.1 is given in §3. We illustrate the result with several corollaries and numerical examples. First, in the following corollary, we show that setting $M=0$ recovers Theorem 1.1 of Steinerberger [23].
Corollary 1.1 ($M=0$).
Additionally, assume $M=0$, then (10) reduces to randomized Kaczmarz (3), and (11) reduces to
(12) $\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=\begin{bmatrix}1-\eta_{l}\\ 0\\ \end{bmatrix}^{\top}\begin{bmatrix}1-\eta_{l}&0\\ -1&\beta\\ \end{bmatrix}^{k}\begin{bmatrix}1\\ -\frac{1}{1-\beta}\end{bmatrix}\langle x_{0}-x,v_{l}\rangle,$
where we used the notation $\eta_{l}:=\sigma_{l}^{2}/\|A\|_{F}^{2}$. Recall that
$\begin{bmatrix}a&0\\ b&c\end{bmatrix}^{n}=\begin{bmatrix}a^{n}&0\\ x_{n}&c^{n}\end{bmatrix},$
where $x_{n}=b(a^{n}+a^{n-1}c+\cdots+ac^{n-1}+c^{n})$. Thus, (12) reduces so
$\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=(1-\eta_{l})^{k+1}\langle x_{0}-x,v_{l}\rangle,$
which is Theorem 1.1 of Steinerberger [23].
Our next corollary optimizes the value of the smoothing parameter $\beta$ by minimizing the maximum magnitude eigenvalue of the $2\times 2$ matrix
(13) $B:=\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}$
from (11) in Theorem 1.1. The eigenvalues of $B$ are
(14) $\lambda_{1}:=\frac{r+\beta+\sqrt{(r-\beta)^{2}-4\zeta}}{2}\quad\text{and}\quad\lambda_{2}:=\frac{r+\beta-\sqrt{(r-\beta)^{2}-4\zeta}}{2}.$
Since $r+\beta>0$, it follows that $|\lambda_{1}|\geq|\lambda_{2}|$ for all $\beta,M\in[0,1]$. In the following, we consider the problem of minimizing $|\lambda_{1}|$. Fix $M\in[0,1]$, and consider the eigenvalue $\lambda_{1}=\lambda_{1}(\beta)$ as a function of the smoothing parameter $\beta$.
Corollary 1.2 (Minimizing $\lambda_{1}$).
Fix $M\in[0,1]$. We have
[…]
The proof of Corollary 1.2 is given in §3.3. Our next corollary restates Theorem 1.1 for the case where $\beta$ is chosen using the optimized value of Corollary 1.2 for the case $0\leq M\leq(1-\sqrt{\eta_{l}})^{2}$, which is the numerical region of interest, see §2.4.
Corollary 1.3.
In addition to the hypotheses of Theorem 1.1 assume that $0\leq M\leq(1-\sqrt{\eta_{l}})^{2}$ and set
(15) $\beta=1-\frac{\eta_{l}}{(1-\sqrt{M})^{2}}.$
Then,
$\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}\left(1+\eta_{l}\frac{\sqrt{M}(k+1))-1}{1-\sqrt{M}}\right)\langle x_{0}-x,v_{l}\rangle.$
The proof of Corollary 1.3 is given in §3.3. Several numerical experiments illustrating Corollary 1.3, and its effectiveness at modeling the behavior of KGSM are included in §2. Before presenting these numerical examples, we emphasize the limitations of our analysis.
Remark 1.1 (Limitations).
We emphasize that Theorem 1.1 and its corollaries provide formulas for the expected value of the signed quantity $\langle x_{k}-x,v_{l}\rangle$. If the sign of this quantity was known to remain constant $\operatorname{sign}\langle x_{k}-x,v_{l}\rangle=\operatorname{sign}\langle x_{0}-x,v_{l}\rangle$ for all $k$, then these results become effective in describing the error $|\langle x_{k+1}-x,v_{l}\rangle|$ of $x_{k}$ to the solution $x$ in direction $v_{l}$. Informally speaking, the sign remaining constant corresponds to $x_{k}$ not overshooting the solution $x$ in the direction $v_{l}$ and keeping a constant approach direction. On the other hand, if $\operatorname{sign}\langle x_{k}-x,v_{l}\rangle$ is not determined by $\operatorname{sign}\langle x_{0}-x,v_{l}\rangle$, then these results may or may not be effective in describing the error $|\langle x_{k+1}-x,v_{l}\rangle|$. For example, the $\operatorname{sign}\langle x_{k}-x,v_{l}\rangle$ could oscillate between positive and negative values or have an equal chance of becoming a large negative or positive value. One possible approach to understanding this issue would be to establish an absolute convergence result or an analog of [23, Theorem 1.3]; see §4 for further discussion.
In the following section, we present various numerical examples that illustrate the dynamics of KGSM, including cases where the dynamics are explained by Theorem 1.1 and cases that raise questions, which we discuss in §4.
2. Numerical Examples
2.1. Notation and preliminaries
Suppose that $A$ is an $m\times n$ matrix with $m\geq n$. Let $\sigma_{1}\geq\cdots\geq\sigma_{n}$ denote the singular vectors of $A$. Using this notation, the Frobenius norm $\|A\|_{F}$ can be expressed by
$\|A\|_{F}=\sqrt{\sum_{j=1}^{n}\sigma_{j}^{2}}.$
In the following, we generate random $m\times n$ matrices with specified singular values $\sigma=(\sigma_{1},\ldots,\sigma_{n})$ as follows. First, we construct a random $m\times n$ matrix $U$ with orthonormal columns by starting with an $m\times n$ matrix $G$ with random independent standard normal entries and applying the Gram-Schmidt procedure to obtain an orthonormal basis for the column space. This construction guarantees that the column space of $U$ is sampled uniformly (with respect to the Haar measure) on the Grassmann manifold ${\bf{Gr}}_{n}(\mathbb{R}^{m})$ of all $n$-dimensional subspaces of $\mathbb{R}^{m}$, see [27, §5.2.6]. Second, using the same procedure, we choose a random $n\times n$ matrix $V$ with orthonormal columns and set
$A:=U\operatorname{diag}(\sigma)V,$
where $\operatorname{diag}(\sigma)$ is the $n\times n$ matrix whose $i$-th diagonal entry is $\sigma_{i}$, and which is zero off the diagonal.
2.2. Basic Example
We generate a matrix with only one small singular value to create a simple example illustrating Corollary 1.3. More precisely, using the procedure described in §2.1, we generate a random $100\times 20$ matrix $A$ whose singular values are
$\sigma_{1}=\cdots=\sigma_{19}=1,\quad\text{and}\quad\sigma_{20}=1/50.$
We choose a random solution vector $x\in\mathbb{R}^{n}$ (with independent standard normal entries) and define $b:=Ax$. We set
(16) $M=0.9,\quad\text{and}\quad\beta=1-\frac{\eta_{20}}{(1-\sqrt{M})^{2}},$
where $\eta_{20}:=\sigma_{20}^{2}/\|A\|_{F}^{2}$, see Remark 2.1 for a discussion about setting the momentum parameter $M$.
We choose a random initial vector $x_{0}$ (with independent standard normal entries) and run randomized Kaczmarz (3) and KGSM (10). At each iterate, we compute the absolute error in the direction of the smallest right singular vector $|\langle x-x_{k},v_{20}\rangle|$, see Figure 2. For comparison, we plot the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Corollary 1.3, respectively.
Figure 2. The numerical error $|\langle x_{k}-x,v_{20}\rangle|$ for randomized Kaczmarz (3) and KGSM (10), and the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Corollary 1.3, for the example of §2.2.
The numerical results in Figure 2 are an example of where the theoretical results of (5) and Corollary 1.3 for the expected signed error $\mathbb{E}\langle x_{k}-x,v_{20}\rangle$ provide an accurate model for the absolute numerical error $|\langle x_{k}-x,v_{20}\rangle|$. Moreover, in this case, the addition of geometrically smoothed momentum drastically increases the convergence rate.
2.3. Complex pertubation
Recall that from (14) that
(17) $\lambda_{1}:=\frac{r+\beta+\sqrt{(r-\beta)^{2}-4\zeta}}{2}$
is the largest magnitude eigenvalue of the $2\times 2$ matrix (13) from Theorem 1.1, where $r=1-\sigma_{l}^{2}/\|A\|_{F}^{2}+M(1-\beta)$ and $\zeta=M(1-\beta)^{2}$. The formula for $\beta$ from Corollary 1.3
$\beta=1-\frac{\eta_{l}}{(1-\sqrt{M})^{2}},$
is the curve where the discriminant $D=(r-\beta)^{2}-4\zeta$ in (17) vanishes. Above the curve, the eigenvalues are complex, while below the curve, there are two distinct real eigenvalues.
In the following, we repeat the same experiment as in §2.2, but perturb $\beta$ by $0.001$ into the region where the eigenvalues are complex; that is, we set
$M=0.9\quad\text{and}\quad\beta=1-\frac{\eta_{20}}{(1-\sqrt{M})^{2}}+0.001.$
In this case, the theoretical estimate for $|\mathbb{E}\langle x-x_{k},v_{20}\rangle|$ for KGSM is computed using Theorem 1.1, and the complex eigenvalues of the matrix in (11) predict oscillatory behavior where the error dips at regular intervals, see Figure 3.
Figure 3. The numerical error $|\langle x_{k}-x,v_{20}\rangle|$ for randomized Kaczmarz (3) and KGSM (10), and the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Theorem 1.1 for the example in §2.3
Informally speaking, Figure 3 illustrates a case where too much momentum is built in a given direction. The following example further explores the parameter space $(M,\beta)$.
2.4. Exploring the $(M,\beta)$ parameter space
This section further explores the $(M,\beta)$ parameter space. For the same linear system described in §2.2, we run KGSM for four different values of $(M,\beta)$; in particular, we consider
(18) $(M,\beta)=(0.85,0.992),(0.9,0.992),\,(0.95,0.992),\text{ and }(0.965,0.932),$
see Figure 4.
Figure 4. Visualization of the values of $(M,\beta)$ from (18). The curve $\beta=1-\eta_{20}/(1-\sqrt{M})^{2}$ is plotted for reference in blue.
For each of the four values of $(M,\beta)$ illustrated in Figure 4, we run KGSM (10), compute the absolute error in the direction of the smallest right singular vector $|\langle x-x_{k},v_{20}\rangle|$, and plot the results in Figure 5.
To interpret the results in Figures 5, recall that the curve $\beta=1-\eta_{20}/(1-\sqrt{M})^{2}$ in Figure 4 divides the $(M,\beta)$ parameter space: above the curve, the eigenvalues of the $2\times 2$ matrix in Theorem 1.1 are complex; below the curve, the eigenvalues are distinct real values; and on the curve, the eigenvalues are repeated real values, see §2.3.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 5. The error $|\langle x_{k}-x,v_{20}\rangle|$ for the randomized Kaczmarz (3) and KGSM (10) for parameters $(M,\beta)$ indicated by markers labeling each plot, which correspond to the markers in Figure 4.
In the following, we briefly discuss the plot corresponding to each marker:
Black circle
The parameters denoted by the black circle marker are those from (16) previously discussed in §2.2 and are included in Figure 5 to provide context.
Red square
The parameters denoted by the red square marker are in the region where the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ are real. The momentum parameter $M$ is smaller compared to the black circle marker, and as a result, the rate of convergence of KGSM is relatively slower; see Figure 5.
Green plus
The parameters denoted by the green plus marker are in the region where the eigenvalues $\lambda_{1}$ and $\lambda_{2}$ are complex, which results in oscillatory behavior; see Figure 5 and the discussion in §2.3.
Blue triangle
The parameters denoted by the blue triangle marker are on the curve $\beta=1-\eta_{20}/(1-\sqrt{M})^{2}$. The momentum parameter $M$ is larger compared to the black circle marker. Here, one might expect that KGSM converges faster relative to the black circle experiment. Instead, the method breaks down, and the iteration diverges; see Figure 5. The following remark discusses setting the momentum parameter $M$.
Remark 2.1 (Setting the momentum parameter $M$).
Recall that by Corollary 1.3, the expected signed error in the direction of the smallest singular vector $\mathbb{E}\langle x-x_{k},v_{20}\rangle$ decreases by a factor of $1-\eta_{l}/(1-\sqrt{M})$ each iteration. We can make this factor arbitrarily small by setting $M$ close to $(1-\sqrt{\eta_{l}})^{2}$. Numerical results indicate that each linear system has a critical value of $M$ beyond which the KGSM iteration fails to converge. Numerically, an effective value of $M$ can be determined by increasing $M$ until the methods diverge or using bisection to identify such a value. We pose several questions related to the critical value of $M$ in §4.
2.5. Linear distribution of singular values
Thus far, we have considered the simplest case, where the matrix $A$ only has one small singular value. Here, we consider cases where the singular values are distributed linearly. In this case, the KGSM method is still effective in accelerating the convergence rate, but the dynamics are no longer precisely predicted by Theorem 1.1.
Using the procedure described in §2.1, we generate a random $100\times 20$ matrix $A$ whose singular values are
$\sigma_{1}=1,\ \sigma_{2}=\frac{19}{20},\quad\cdots\quad,\sigma_{20}=\frac{1}{20}.$
We set
$M=0.85,\quad\text{and}\quad\beta=1-\frac{\eta_{20}}{(1-\sqrt{M})^{2}}.$
We choose a random initial vector $x_{0}$ (with independent standard normal entries) and run randomized Kaczmarz (3) and KGSM (10). At each iterate, we compute the absolute error in the direction of the smallest right singular vector $|\langle x-x_{k},v_{20}\rangle|$, see Figure 6. For comparison, we plot the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Corollary 1.3, respectively.
Figure 6. The numerical error $|\langle x_{k}-x,v_{20}\rangle|$ for randomized Kaczmarz (3) and KGSM (10), and the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Corollary 1.3, for the example of §2.5.
The numerical example in Figure 6 has two notable attributes: first, even though the parameters are set so they are on the curve $\beta=1-\frac{\eta_{20}}{(1-\sqrt{M})^{2}}$, the error dips down at the beginning as if the eigenvalues were slightly complex. We attribute this dip to numerical instability. Second, while the numerical error is no longer precisely predicted by Corollary 1.3, it appears the estimated rate approaches the predicted rate over time.
2.6. Many small singular values
The numerical results of the previous section indicate that the dynamics of KGSM are more complicated when the underlying matrix has many small singular values. In the following, we consider the extreme case of a matrix where all of the singular values are small except for one. We repeat the same numerical experiment (with the parameters) described in §2.5 but construct the matrix $A$ to have singular values
$\sigma_{1}=1,\quad\text{and}\quad\sigma_{2}=\cdots=\sigma_{20}=1/50.$
The results are plotted in Figure 7.
Figure 7. The numerical error $|\langle x_{k}-x,v_{20}\rangle|$ for randomized Kaczmarz (3) and KGSM (10), and the theoretical estimates for $|\mathbb{E}\langle x_{k}-x,v_{20}\rangle|$ from (5) and Corollary 1.3, for the example of §2.6.
The numerical example in Figure 7 exhibits complicated dynamics. Observe that the theoretical estimate for the signed expected error from Corollary 1.3 no longer seems to match the numerical error $|\langle x-x_{k},v_{20}\rangle|$, even for a large number of iterations. Furthermore, the error includes multiple dips similar to the complex dynamics of §2.3, but without a defined period. This example was designed as a worst-case situation for KGSM.
2.7. Comparison to $\ell^{2}$-norm error
So far, we have been plotting the error $|\langle x_{k}-x,v_{n}\rangle|$ in the direction of the smallest singular value $v_{n}$. In this section, we compare the singular vector error $|\langle x_{k}-x,v_{n}\rangle|$ to the $\ell^{2}$-norm error $\|x-x_{k}\|_{2}$. We start by considering the numerical example described in §2.3 and provide a side-by-side comparison of error in the direction of the smallest singular vector and the $\ell^{2}$-norm error in Figure 8.
Refer to caption Refer to caption
Figure 8. The error $|\langle x_{k}-x,v_{20}\rangle|$ (left) and $\|x_{k}-x\|_{2}$ (right) for randomized Kaczmarz (3) and KGSM (10) for a linear system with one small singular vector.
In the numerical example of Figure 8, where the underlying matrix only has one small singular value, the square error is roughly a scaled version of the error in the direction of the smallest singular vector. Next, we modify this numerical example by adding another small singular value. More precisely, we repeat the same numerical experiment with the same parameters but now set:
$\sigma_{1}=\cdots=\sigma_{18}=1,\quad\text{and}\quad\sigma_{19}=\sigma_{20}=1/50.$
We plot the error in the direction of one of the smallest singular vectors in comparison to the $\ell^{2}$-norm error in Figure 9.
Refer to caption Refer to caption
Figure 9. The error $|\langle x_{k}-x,v_{20}\rangle|$ (left) and $\|x_{k}-x\|_{2}$ (right) for randomized Kaczmarz (3) and KGSM (10) for a linear system with two small singular vectors.
In the numerical example in Figure 9, the $\ell^{2}$-norm error does not capture the finer features of the singular vector error since there are two small singular vectors whose errors are not perfectly synchronized due to numerical error. The general rate of decrease is similar, but the finer features are lost.
3. Proof of main result
This section gives the proof of Theorem 1.1 and its corollaries.
3.1. Notation and preliminaries
Let $A$ be $m\times n$ matrix that is tall in the sense that $m\geq n$. Let $v_{l}$ denote the right singular vector of $A$ associated with the $l$-th largest singular value $\sigma_{l}$. The proof of the results uses the following lemma, which is the first step in the proof of Steinerberger [23, Theorem 1.1].
Lemma 3.1.
Suppose that $a^{\top}$ is a row of $A$ chosen with probability $\|a\|_{2}^{2}/\|A\|_{F}^{2}$. Then,
$\mathbb{E}\frac{\langle x,a\rangle\langle a,v_{l}\rangle}{\|a\|_{2}^{2}}=\frac{\sigma_{l}^{2}}{\|A\|_{F}^{2}}\langle x,v_{l}\rangle,$
for any $x\in\mathbb{R}^{n}$.
We give the proof for completeness.
Proof of Lemma 3.1.
Direct calculation gives
$\mathbb{E}\frac{\langle x,a\rangle\langle a,v_{l}\rangle}{\|a\|_{2}^{2}}=\sum_{i=1}^{m}\frac{\|a_{i}\|_{2}^{2}}{\|A\|_{F}^{2}}\frac{x^{\top}a_{i}a_{i}^{\top}v_{l}}{\|a_{i}\|_{2}^{2}}=\frac{x^{\top}A^{\top}Av_{l}}{\|A\|_{F}^{2}}=\frac{\sigma_{l}^{2}x^{\top}v_{l}}{\|A\|_{F}^{2}},$
which completes the proof. ∎
3.2. Proof of main result
Proof of Theorem 1.1.
Given an initial $x_{0}$ and $\beta\in[0,1)$ and $M\in[0,1]$ let $y_{0}=0\in\mathbb{R}^{n}$ be the zero vector. Suppose that $x_{k}$ and $y_{k}$ are defined iteratively by (10). Without loss of generality, by the possibility of replacing $x_{k}$ with $x_{k}-x$, we can assume the solution $x$ is the zero vector. In this case, (10) reduces to
(19) $\left\{\begin{array}[]{ccl}x_{k+1}&=&x_{k}-\displaystyle\frac{\langle x_{k},a_{i_{k}}\rangle a_{i_{k}}}{\|a_{i_{k}}\|^{2}_{2}}+My_{k}\\[10.0pt] y_{k+1}&=&\beta y_{k}+(1-\beta)(x_{k+1}-x_{k}).\end{array}\right.$
Inducting on $k$ gives
(20) $y_{k}=(1-\beta)\sum_{j=1}^{k}\beta^{k-j}(x_{j}-x_{j-1}).$
Indeed, the base case $k=0$ of the induction holds since an empty sum is equal to zero, and assuming the formula holds for $y_{k}$ and applying (19) gives
(21) $\begin{split}y_{k+1}&=\beta\left((1-\beta)\sum_{j=1}^{k}\beta^{k-j}(x_{j}-x_{j-1})\right)+(1-\beta)(x_{k+1}-x_{k}),\\ &=(1-\beta)\sum_{j=1}^{k+1}\beta^{k+1-j}(x_{j}-x_{j-1}).\end{split}$
Observe that splitting the sum over $j$ in (20) into two sums, shifting the indices, and combining terms gives
(22) $\begin{split}\sum_{j=1}^{k}\beta^{k-j}(x_{j}-x_{j-1})&=\sum_{j=1}^{k}\beta^{k-j}x_{j}-\sum_{j=1}^{k}\beta^{k-j}x_{j-1},\\ &=\sum_{j=1}^{k}\beta^{k-j}x_{j}-\sum_{j=0}^{k-1}\beta^{k-j-1}x_{j},\\ &=x_{k}-(1-\beta)\left(\sum_{j=1}^{k-1}\beta^{k-j-1}x_{j}\right)-\beta^{k-1}x_{0}.\end{split}$
It follows that
(23) $y_{k}=(1-\beta)x_{k}-(1-\beta)^{2}\left(\sum_{j=1}^{k-1}\beta^{k-j-1}x_{j}\right)-(1-\beta)\beta^{k-1}x_{0}.$
Combining (19), (20), (21), and (22) gives
(24) $\begin{split}x_{k+1}=&\big{(}1+M(1-\beta)\big{)}x_{k}-\frac{\langle x_{k},a_{i_{k}}\rangle}{\|a_{i_{k}}\|_{2}^{2}}a_{i_{k}}\\ &-M(1-\beta)^{2}\left(\sum_{j=1}^{k-1}\beta^{k-j-1}x_{j}\right)-M(1-\beta)\beta^{k-1}x_{0}.\end{split}$
Recall that $v_{l}$ denotes the right singular vector associated with the $l$-th largest singular value $\sigma_{l}$ of $A$. Taking the inner product of (24) with $v_{l}$ gives
$\begin{split}\langle x_{k+1},v_{l}\rangle=&\big{(}1+M(1-\beta)\big{)}\langle x_{k},v_{l}\rangle-\frac{\langle x_{k},a_{i_{k}}\rangle\langle a_{i_{k}},v_{l}\rangle}{\|a_{i_{k}}\|^{2}_{2}}\\ &-M(1-\beta)^{2}\left(\sum_{j=1}^{k-1}\beta^{k-j-1}\langle x_{j},v_{l}\rangle\right)-M(1-\beta)\beta^{k-1}\langle x_{0},v_{l}\rangle.\end{split}$
Let $\mathbb{E}_{i_{0},\ldots,i_{k}-1}$ denote the expectation conditional on $i_{0},\ldots,i_{k-1}$. We have
(25) $\begin{split}\mathbb{E}_{i_{0},...,i_{k-1}}\langle x_{k+1},v_{l}\rangle=&\langle x_{k},v_{l}\rangle\left(1-\frac{\sigma_{l}^{2}}{\|A\|_{F}^{2}}+M(1-\beta)\right)\\ &-M(1-\beta)^{2}\left(\sum_{j=1}^{k-1}\beta^{k-j-1}{\langle x_{j},v_{l}\rangle}\right)-M(1-\beta)\beta^{k-1}\langle x_{0},v_{l}\rangle,\end{split}$
where the term $\sigma_{l}^{2}/\|A\|_{F}^{2}$ results from applying Lemma 3.1. Set
(26) $\begin{split}r&:=1-\frac{\sigma_{l}^{2}}{\|A\|_{F}^{2}}+M(1-\beta),\\ \zeta&:=M(1-\beta)^{2},\\ S_{n}&:=\left(\sum_{j=1}^{n}\beta^{n-j}{\langle x_{j},v_{l}\rangle}\right)+\frac{\beta^{n}}{1-\beta}\langle x_{0},v_{l}\rangle.\end{split}$
With this notation, (25) can be written as
(27) $\mathbb{E}_{i_{0},...,i_{k-1}}\langle x_{k+1},v_{l}\rangle=r\langle x_{k},v_{l}\rangle-\zeta S_{k-1}.$
For any fixed integer $k\geq 1$, and $j=1,2,\ldots,k$, we claim that
(28) $\mathbb{E}_{i_{0},...,i_{k-j}}\langle x_{k+1},v_{l}\rangle=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{j-1}\begin{bmatrix}\langle x_{k+1-j},v_{l}\rangle\\ -S_{k-j}\end{bmatrix}.$
For fixed integer $k\geq 1$, we prove this expression by induction on $j$. In base case $j=1$ of the induction, (28) reduces to (27). Assume that (28) holds up to $j$. By the tower property of conditional expectation, we have $\mathbb{E}_{i_{0},...,i_{k-j-1}}\mathbb{E}_{i_{0},...,i_{k-j}}=\mathbb{E}_{i_{0},...,i_{k-j-1}}$. Thus, taking the conditional expectation $\mathbb{E}_{i_{0},...,i_{k-j-1}}$ of both sides gives
$\begin{split}\mathbb{E}_{i_{0},...,i_{k-j-1}}\langle x_{k+1},v_{l}\rangle&=\mathbb{E}_{i_{0},...,i_{k-j-1}}\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{j-1}\begin{bmatrix}\langle x_{k-j+1},v_{l}\rangle\\ -S_{k-j}\end{bmatrix},\\ &=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{j-1}\mathbb{E}_{i_{0},...,i_{k-j-1}}\begin{bmatrix}\langle x_{k-j+1},v_{l}\rangle\\ -S_{k-j}\end{bmatrix},\\ &=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{j-1}\begin{bmatrix}r\langle x_{k-j},v_{l}\rangle-\zeta S_{k-j-1}\\ -\langle x_{k-j},v_{l}\rangle-\beta S_{k-j-1}\end{bmatrix},\\ &=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{j}\begin{bmatrix}\langle x_{k-j},v_{l}\rangle\\ -S_{k-j-1}\end{bmatrix},\\ \end{split}$
where the second equality follows from the linearity of expectation, the third equality follows from (27) and the definition of $S_{k}$, and the fourth equality follows from factoring out
$\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}$
from the right vector. Setting $j=k-1$ gives
(29) $\mathbb{E}_{i_{0}}\langle x_{k+1},v_{l}\rangle=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k-1}\begin{bmatrix}\langle x_{1},v_{l}\rangle\\ -S_{0}\end{bmatrix}.$
Taking the full expectation of (29) gives
(30) $\begin{split}\mathbb{E}\langle x_{k+1},v_{l}\rangle&=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k-1}\mathbb{E}\begin{bmatrix}\langle x_{1},v_{l}\rangle\\ -S_{0}\end{bmatrix},\\ &=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k-1}\begin{bmatrix}(1-\sigma_{l}^{2}/\|A\|_{F}^{2})\langle x_{0},v_{l}\rangle\\ -\frac{1}{1-\beta}\langle x_{0},v_{l}\rangle\end{bmatrix},\\ &=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k-1}\begin{bmatrix}r-\frac{\zeta}{1-\beta}\\ -\frac{1}{1-\beta}\end{bmatrix}\langle x_{0},v_{l}\rangle,\end{split}$
where the final equality follows from the definition of $r$ and $\zeta$, see (26). In addition since
$\begin{bmatrix}r&\zeta\\ -1&\beta\end{bmatrix}\begin{bmatrix}1\\ \frac{-1}{1-\beta}\end{bmatrix}=\begin{bmatrix}r-\frac{\zeta}{1-\beta}\\ -\frac{1}{1-\beta}\end{bmatrix},$
we have proved that when $k\geq 1$, we have
(31) $\mathbb{E}\langle x_{k+1},v_{l}\rangle=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k}\begin{bmatrix}1\\ -\frac{1}{1-\beta}\end{bmatrix}\langle x_{0},v_{l}\rangle.$
Note that so far, we have assumed $k\geq 1$. To complete the proof, it remains to consider the case when $k=0$. On one hand computing the expectation $\mathbb{E}\langle x_{1},v_{l}\rangle$ using (19) gives
(32) $\mathbb{E}\langle x_{1},v_{l}\rangle=\langle x_{0},v_{l}\rangle\left(1-\frac{\sigma_{l}^{2}}{\|A\|_{F}^{2}}\right)=\left(r-\frac{\zeta}{1-\beta}\right)\langle x_{0},v_{l}\rangle;$
on the other hand, setting $k=0$ on the right-hand side of (31) gives
(33) $\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}1\\ -\frac{1}{1-\beta}\end{bmatrix}\langle x_{0},v_{l}\rangle=\left(r-\frac{\zeta}{1-\beta}\right)\langle x_{0},v_{l}\rangle.$
Since the right-hand side of (32) and (33) are equal, we conclude that the case $k=0$ holds, and thus we have established
(34) $\mathbb{E}\langle x_{k+1},v_{l}\rangle=\begin{bmatrix}r\\ \zeta\\ \end{bmatrix}^{\top}\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}^{k}\begin{bmatrix}1\\ -\frac{1}{1-\beta}\end{bmatrix}\langle x_{0},v_{l}\rangle,$
for all $k\geq 0$. This completes the proof. ∎
3.3. Proof of corollaries
Proof of Corollary 1.2.
Let $M\in[0,1]$ and $0<\eta_{l}\leq 1$ be fixed. By (14) we have
(35) $\lambda_{1}(\beta):=\frac{r+\beta+\sqrt{(r-\beta)^{2}-4\zeta}}{2},$
where
$r=1-\eta_{l}+M(1-\beta),\quad\zeta=M(1-\beta)^{2},\quad\text{and}\quad\eta_{l}=\frac{\sigma_{l}^{2}}{\|A\|_{F}^{2}}.$
The discriminant $D:=(r-\beta)^{2}-4\zeta$ has two zeros
$\beta_{0}:=1-\frac{\eta_{l}}{(1-\sqrt{M})^{2}},\quad\text{and}\quad\beta_{1}:=1-\frac{\eta_{l}}{(1+\sqrt{M})^{2}}.$
The discriminant $D$ is nonnegative when $\beta\in(-\infty,\beta_{0}]$, negative when $\beta\in(\beta_{0},\beta_{1})$, and nonnegative when $\beta\in[\beta_{1},\infty)$.
Case 1: $M\in[0,(1-\sqrt{\eta_{l}})^{2}]$
In this case, $\beta_{0}\geq 0$ and we claim
$\beta_{0}=\operatorname*{argmin}_{\beta\in[0,1]}|\lambda_{1}(\beta)|.$
We prove this claim by considering three subcases.
Case 1a: $\beta\in[0,\beta_{0}]$
In this case, the discriminant $D$ is nonnegative so
$|\lambda_{1}(\beta)|=\frac{1}{2}\left(\beta+r+\sqrt{(\beta+r)^{2}-4(r\beta+\zeta)}\right)=\frac{1}{2}\left(\beta+r+\sqrt{D}\right).$
We claim that $\partial_{\beta}|\lambda_{1}(\beta)|\leq 0$. Indeed,
$\partial_{\beta}|\lambda_{1}(\beta)|=\frac{1}{2}\left(1-M+\frac{\partial_{\beta}D}{2\sqrt{D}}\right),$
where
$\frac{\partial_{\beta}D}{2}=-(1-\beta)(1-M)^{2}+\eta_{l}(1+M).$
The assumption that $\beta\in[0,\beta_{0}]$ and $M\in[0,(1-\sqrt{\eta_{l}})^{2}]$ implies that $\partial_{\beta}D/2\leq 0$. Thus, to show that $\partial_{\beta}|\lambda_{1}(\beta)|\leq 0$ it suffices to show that
$(1-M)^{2}\leq\left(\frac{\partial_{\beta}D}{2\sqrt{D}}\right)^{2}.$
That is, it suffices to show that
$(1-M)^{2}D-\left(\frac{\partial_{\beta}D}{2}\right)^{2}\leq 0.$
Using the definition of $D$ and expanding terms gives
$(1-M)^{2}D-\left(\frac{\partial_{\beta}D}{2}\right)^{2}=-4\eta_{l}^{2}M\leq 0,$
so we conclude that $\partial_{\beta}|\lambda_{1}(\beta)|\leq 0$, which implies that $|\lambda_{1}(\beta)|\geq|\lambda_{1}(\beta_{0})|$ for $\beta\in[0,\beta_{0}]$.
Case 1b: $\beta\in(\beta_{0},\beta_{1})$
In this case, the discriminant is negative so
$|\lambda_{1}(\beta)|^{2}=\frac{(r+\beta)+4\zeta-(r-\beta)^{2}}{4}=(1-\eta_{l}+M(1-\beta))\beta+M(1-\beta)^{2}.$
Observe that the derivative of $|\lambda_{1}(\beta)|^{2}$ with respect to $\beta_{0}$ is positive
$\partial_{\beta}|\lambda_{1}(\beta)|^{2}=1-\eta_{l}-M\geq 1-\eta_{l}-(1-\sqrt{\eta_{l}})^{2}=2(\sqrt{\eta_{l}}-\eta_{l})>0,$
which implies $|\lambda_{1}(\beta_{0})|<|\lambda_{1}(\beta)|$ for all $\beta\in(\beta_{0},\beta_{1})$.
Case 1c: $\beta\in[\beta_{1},1)$
In this case, $r\leq\beta$, so we can estimate
$|\lambda_{1}(\beta)|\geq\frac{r+\beta}{2}\geq r\geq 1-\eta_{l}\geq|\lambda(\beta_{0})|,$
which completes the proof of Case 1.
Case 2: $M\in((1-\sqrt{\eta_{l}})^{2},1)$
In this case, we claim
$\operatorname*{argmin}_{\beta\in[0,1]}{|\lambda_{1}(\beta)|}=\left\{\begin{array}[]{cl}0&\text{if }M\in[(1-\sqrt{\eta_{l}})^{2},1-\eta_{l}]\\ \beta_{1}&\text{if }M\in(1-\eta_{l},1]\end{array}\right..$
We prove the claim by considering two subcases.
Case 2a : $\beta\in[0,\beta_{1})$
Since the assumption on $M$ implies $[0,\beta_{1})\subset(\beta_{0},\beta_{1})$ the discriminant is negative and
$|\lambda_{1}(\beta)|^{2}=\frac{(\beta+r)^{2}+4(r\beta+\zeta)-(\beta+r)^{2}}{4}=(1-\eta_{l}+M(1-\beta))\beta+M(1-\beta)^{2}.$
Taking the derivative with respect to $\beta$ gives
$\partial_{\beta}|\lambda_{1}(\beta)|^{2}=1-\eta_{l}-M.$
We conclude that the minimum occurs at $\beta=0$ when $M\in[(1-\sqrt{\eta_{l}})^{2},1-\eta]$ or at $\beta_{1}$ when $M\in[1-\eta_{l},1]$.
Case 2b: $\beta\in[\beta_{1},1)$
Since $\beta\in[\beta_{1},1)$ the discriminant is nonnegative, the eigenvalues are real, and
$\partial_{\beta}|\lambda_{1}(\beta)|=\frac{1}{2}\left(1-M+\frac{\partial_{\beta}D}{2\sqrt{D}}\right)$
as in Case 1a. Observe that, by our assumption on $\beta$,
$\partial_{\beta}D/2=-(1-\beta)(1-M)^{2}+\eta_{l}(1+M)>\left(\frac{\eta_{l}}{(1+\sqrt{M})^{2}}\right)(1-M)^{2}+\eta_{l}(1+M)$
$=\eta_{l}\left[\frac{1-M^{2}}{(1+\sqrt{M})^{2}}+(1+M)\right].$
Since $M\in((1-\sqrt{\eta_{l}})^{2},1)$, all terms above are positive, which shows $\partial_{\beta}D/2>0$. Furthermore, $\partial_{\beta}|\lambda_{1}(\beta)|>0$ implying the minimum occurs at $\beta_{1}$. This means that for $M\in[(1-\sqrt{\eta})^{2},1]$ the minimum reduces to Case 2a by continuity of $|\lambda_{1}(\beta)|$, completing our proof. ∎
Proof of Corollary 1.3.
Let $B$ denote the $2\times 2$ matrix
$B=\begin{bmatrix}r&\zeta\\ -1&\beta\\ \end{bmatrix}$
from (11) in Theorem 1.1 and recall the eigenvalues of $B$ are
$\lambda_{1}:=\frac{r+\beta+\sqrt{(r-\beta)^{2}-4\zeta}}{2}\quad\text{and}\quad\lambda_{2}:=\frac{r+\beta-\sqrt{(r-\beta)^{2}-4\zeta}}{2}.$
Set
$\beta=1-\frac{\eta_{l}}{(1-\sqrt{M})^{2}}.$
In this case
$B=\begin{bmatrix}1-\eta_{l}\frac{1-2\sqrt{M}}{(1-\sqrt{M})^{2}}&\eta_{l}^{2}\frac{M}{(1-\sqrt{M})^{4}}\\[5.0pt] -1&\eta_{l}\frac{M}{(1-\sqrt{M})^{2}}\end{bmatrix}$
has the eigenvalue
$\lambda=1-\frac{\eta_{l}}{1-\sqrt{M}}$
of algebraic multiplicity two. It is straightforward to verify that the Jordan decomposition of this matrix is
$B=\begin{bmatrix}\frac{\eta_{l}^{2}M}{(1-\sqrt{M})^{4}}&0\\[5.0pt] -\frac{\eta_{l}\sqrt{M}}{(1-\sqrt{M})^{2}}&1\end{bmatrix}\begin{bmatrix}1-\frac{\eta_{l}}{1-\sqrt{M}}&1\\[5.0pt] 0&1-\frac{\eta_{l}}{1-\sqrt{M}}\end{bmatrix}\begin{bmatrix}\frac{(1-\sqrt{M})^{4}}{\eta_{l}^{2}M}&0\\[5.0pt] \frac{(1-\sqrt{M})^{2}}{\eta_{l}\sqrt{M}}&1\end{bmatrix}.$
It follows that
$B^{k}=\begin{bmatrix}\frac{\eta_{l}^{2}M}{(1-\sqrt{M})^{4}}&0\\[5.0pt] -\frac{\eta_{l}\sqrt{M}}{(1-\sqrt{M})^{2}}&1\end{bmatrix}\begin{bmatrix}\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}&k\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k-1}\\[5.0pt] 0&\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}\end{bmatrix}\begin{bmatrix}\frac{(1-\sqrt{M})^{4}}{\eta_{l}^{2}M}&0\\[5.0pt] \frac{(1-\sqrt{M})^{2}}{\eta_{l}\sqrt{M}}&1\end{bmatrix}.$
By Theorem 1.1 we have
$\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=\begin{bmatrix}1-\eta_{l}+M\frac{\eta_{l}}{(1-\sqrt{M})^{2}}\\[5.0pt] M\frac{\eta_{l}^{2}}{(1-\sqrt{M})^{4}}\end{bmatrix}^{\top}\begin{bmatrix}\frac{\eta_{l}^{2}M}{(1-\sqrt{M})^{4}}&0\\[5.0pt] -\frac{\eta_{l}\sqrt{M}}{(1-\sqrt{M})^{2}}&1\end{bmatrix}\\ \begin{bmatrix}\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}&k\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k-1}\\[5.0pt] 0&\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}\end{bmatrix}\begin{bmatrix}\frac{(1-\sqrt{M})^{4}}{\eta_{l}^{2}M}&0\\[5.0pt] \frac{(1-\sqrt{M})^{2}}{\eta_{l}\sqrt{M}}&1\end{bmatrix}\begin{bmatrix}1\\[5.0pt] \frac{-(1-\sqrt{M})^{2}}{\eta_{l}}\end{bmatrix}\langle x_{0}-x,v_{l}\rangle;$
performing matrix multiplication gives
$\mathbb{E}\langle x_{k+1}-x,v_{l}\rangle=\left(1-\frac{\eta_{l}}{1-\sqrt{M}}\right)^{k}\left(1+\frac{\eta_{l}\left((\sqrt{M}(k+1))-1\right)}{1-\sqrt{M}}\right)\langle x_{0}-x,v_{l}\rangle,$
which completes the proof. ∎
4. Discussion
4.1. Summary
This paper introduced the randomized Kaczmarz with geometrically smoothed momentum (KGSM) algorithm. The method depends on two hyperparameters
  • the momentum parameter $M$, and
  • the geometric smoothing parameter $\beta$.
We analyzed the behavior of this method across different distributions of singular vectors and different values of $M$ and $\beta$. Our main result, Theorem 1.1, provides a formula for the expected signed error $\mathbb{E}\langle x-x_{k},v_{l}\rangle$ of KGSM in the direction of the singular vectors $v_{l}$ of the matrix $A$ that defines the least squares loss. In Corollary 1.1, we observed how our result extends Theorem 1.1 of Steinerberger [23], while Corollaries 1.2 and 1.3 optimize the smoothing parameter $\beta$ for a fixed momentum parameter $M$. Our theoretical results are illustrated by several numerical examples in §2. In addition, we provided examples that show the limitations of our analysis and pose questions regarding the dynamics of KGSM. We detail some of these questions below and discuss potential avenues to extend our analysis.
4.2. Limitations and questions
The main limitation of our theoretical analysis is that it does not provide insight into the value of the critical momentum parameter $M$ for a given linear system. We discuss this issue and related questions in the following.
4.2.1. Critical momentum
Our numerical results indicate that there is a critical momentum along the curve illustrated in Figure 4 where the algorithm’s behavior breaks down. More generally, there is a region of the $(M,\beta)$ parameter space where the method fails to converge. What is the critical $M$ or region in $(M,\beta)$ parameter space where this failure occurs?
4.2.2. Estimating $M$ and $\beta$ from data
We did not investigate methods for estimating effective parameters $M$ and $\beta$ from data. Even if formulas for $M$ and $\beta$ are known, these would need to be estimated from data to implement the method for practical problems. Is it possible to estimate suitable parameters as the algorithm runs adaptively?
4.2.3. Direction of convergence and convergence in $\ell^{2}$-norm
We did not establish an analog of [23, Theorem 1.3], which would help to explain any directional change in KGSM. Can this result be generalized to the setting with geometrically smoothed momentum? In particular, can the change in direction be controlled as a function of $M$? Another related question is establishing convergence for the expected absolute error in the direction of a singular vector $|\langle x-x_{k},v_{l}\rangle|$ or for the $\ell^{2}$-norm $\|x-x_{k}\|_{2}$. By Jensen’s inequality, we have $|\mathbb{E}\langle x-x_{k},v_{l}\rangle|\leq\mathbb{E}|\langle x-x_{k},v_{l}\rangle|$, so our results provide a lower bound that our numerical results indicate is sharp in some cases. Thus, at least for some special cases, proving a result about the absolute error may be possible.
4.2.4. Block methods
Our analysis considered the case of block size one (or equivalent minibatch size). Can the analysis be extended to Kaczmarz methods with variable block or minibatch size? One potential approach would be to combine our analysis with that of Bollapragada, Chen, and Ward [1] or to directly approach the problem using the methodology of other block methods.
Acknowledgments
NFM was supported in part by a start-up grant from Oregon State University.